OpenStudy (anonymous):

For f(x) = -x -7, find f(-5)

OpenStudy (y1066380):

Just put x=-5 into -x -7

OpenStudy (y1066380):

Do you get it?

OpenStudy (anonymous):

no i dont

OpenStudy (y1066380):

F(x)= -x -7 F(-5) = -(-5)-7 =5-7 =-2

OpenStudy (anonymous):

ohh okay thank you!!

OpenStudy (y1066380):

You're welcome~